1.  a beam LO weighing 50N rests against a wall, the lower part of the beam extending 15 m from the base of the wall. Calculate the minimum value of the force F that can be applied to keep the beam in equilibrium. The coefficient of friction at L and O is 0.30. Height from the bottom of the ground up to the top of the ladder is 30m. F is not at the middle of the ladder. The componenets acting on the top the ladder are RL and NL. The componenets acting on the top the ladder are R0 and N0. The distance of force F from the ground is 10m.
1
Expert's answer
2021-01-28T12:19:18-0500

Answer

Answer

Let's assume that beam is making angle 45° with horizontal then moment about point O

"50(\frac{15}{2}) \sqrt{2}sin45°+0.3N_2(15) \sqrt{2}cos45°=N_2(15)"

Force required to keep the beam in equilibrium


"35.7(0.3) +P=50cos45°\\P=24.6N"